OpenStudy (anonymous):

Find the limit as x approaches pi of ( (x-pi) / (x+pi) )^1/5

OpenStudy (anonymous):

\[\lim_{x \rightarrow \pi} \sqrt[5]{(x-\pi)\div(x+\pi)}\]

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

I'm sorry to keep bugging everyone, but I'm unable to wrap my mind around this problem? Does anyone know how to do it?